# The continuum limit of some products of random matrices associated with renewing flows

### v1

@@ 7. Discussion and concluding remarks
For $d=2$, we obtained exact formulae for the first two cumulants in terms of elliptic functions with modulus $k=1/\sqrt{2}$.
For $d\geq 2$, we embedded the problem in a $k$-dependent family such that, when $k=0$, the problem is solvable. Treating $k$ as a perturbation parameter, we then obtained expansions in ascending powers of $k^{2}$ for a function $L(k,\ell)$ that coincides with $L(\ell)$ when $k^{2}=1/d$. The first few terms of this expansion were computed explicitly for $d\in\{2,3\}$. We observed that the radius of convergence decreases with $\ell$ so that, when $k^{2}=1/d$, they only converge for moderate values of $\ell$. Nevertheless, they provided an effective means of computing the cumulants, which depend only on the local behaviour of the generalised Lyapunov exponent at $\ell=0$.
We also showed that, for every $0\leq k^{2}\leq 1/d$, $\ell\mapsto L(k,\ell)$ is— still in the continuum limit— the generalised Lyapunov exponent of a product of matrices associated with the renewing flow obtained by inserting time intervals of pure strain; the number $1/d-k^{2}$ then measures the strength of the disorder in the pure strain, relative to that in the pure-strain-free version.
@@ 7.1. Comparison with the calculations of Haynes and Vanneste
In the work of Haynes and Vanneste [10], it is the value taken by $L$ at its minimum, as well as the value of the second derivative there, that are of interest since, under the conditions stated in the introduction, they characterise the variance decay rate of the passive scalar, in the limit of vanishing diffusion. The minimum is achieved when
(7.1) $\ell=\ell_{\min}:=-\frac{d}{2}\,.$
It turns out that $\ell_{\min}$ is small enough to ensure that the expansion of $L(k,\ell)$ in powers of $k^{2}$ approximates $L(\ell)$ well in its vicinity when $k^{2}=1/d$. In particular, when $d=2$, we find
(7.2) $\frac{1}{\tau^{2}}L\left(k,\ell_{\min}\right)=-\frac{1}{2}+\frac{1}{2}k^{2}+\frac{9}{128}k^{4}+\frac{9}{256}k^{6}+\frac{47241}{2097152}k^{8}\\ +\frac{67995}{4194304}k^{10}+\frac{13407669}{1073741824}k^{12}+\frac{21598857}{2147483648}k^{14}+\cdots$
and
(7.3) $\frac{1}{\tau^{2}}\frac{\partial^{2}L}{\partial\ell^{2}}\left(k,\ell_{\min}\right)=1-k^{2}-\frac{5}{32}k^{4}-\frac{5}{64}k^{6}-\frac{13109}{262144}k^{8}\\ -\frac{18847}{524288}k^{10}-\frac{7424679}{268435456}k^{12}-\frac{11948355}{536870912}k^{14}+\cdots$
For $k^{2}=1/2$, this yields
$\frac{1}{\tau^{2}}L\left(\ell_{\min}\right)=-0.2257817708\ldots\;\;\text{and}\;\;\frac{1}{\tau^{2}}L^{\prime\prime}\left(\ell_{\min}\right)=0.446190238842\ldots$
This agrees with the numerical values reported by Haynes and Vanneste [10] in their §III, but the correspondence is clouded by some unfortunate clashes in the notation. To clarify the correspondence, let us take as starting point Haynes & Vanneste’s Equation (3.19):
@@
$\sigma\sim\ell+1\,,\;\;f_{\sigma}(\theta)\sim v(\theta)\,,\;\;\lambda\sim\frac{1}{\tau^{2}}\left(1-\lambda\right)\,,\;\;\Lambda(\sigma)\sim\frac{1}{\tau^{2}}\left[1-\lambda(\ell)\right]\,.$
Recall that, in our notation,
$\lambda(\ell)=\exp\left[L(\ell)\right]\,.$
In the continuum limit in which we systematically neglect the terms of $o(\tau^{2})$ as $\tau\rightarrow 0+$, one therefore ends up with the correspondence
$\Lambda(\sigma)\sim-\frac{1}{\tau^{2}}L(\ell)\,,\;\;\sigma\sim\ell+1\,.$
In the statement just above their equation (3.21), Haynes and Vanneste write $\Lambda^{\prime\prime}(0)\approx 0.45$, but when one compares their equations (3.8) and (3.21), it becomes clear that $\Lambda^{\prime\prime}(0)\approx-0.45$ was intended.
@@ 7.3. Beyond the continuum limit
where, according to Equation (1.5), $\alpha=2\tau\cos\theta$ and $\beta=2\tau\cos\theta$, with $\theta$ uniformly distributed in $(-\pi,\pi)$.
Let us take the calculation of the first cumulant as a modest objective. The following result, whose proof we omit, is a simple application of Furstenberg’s formula [8] for the Lyapunov exponent of a product of independent identically-distributed random matrices:
@@
${\gamma}_{1}=\int_{\mathbb{R}}\nu(dz)\,{\mathbb{E}}\left(\ln\left|\frac{\alpha+z}{z}\right|\right)\,.$
@@
${\gamma}_{1}=-2\int_{\mathbb{R}}\nu(dz)\,\ln|z|\,.$
The problem of computing $\gamma_{1}$ therefore reduces to that of finding the distribution of the random continued fraction implied by Equation (7.7). This continued fraction has the same form as that considered by Dyson in his study of a random linear chain [6, 7]. There is, however, one significant novelty: in Dyson’s case the $\alpha_{j}$ and $\beta_{j}$ represent random masses and Hooke constants, and so assume strictly positive values; by contrast, in the present case, the $\alpha_{j}$ and $\beta_{j}$ can assume negative values. This is problematic (and the situation gets worse as $\tau$ increases) because Equation (7.7) implies that the probability of dividing by a number in an interval centered on zero is proportional to the size of the interval.
@@
${\gamma}_{1,0}=0\,,\;\;{\gamma}_{1,1}=\zeta\,,\;\;{\gamma}_{1,2}=\frac{1}{24}+\frac{3}{8}\zeta^{2}\,,\;\;{\gamma}_{1,3}=\frac{5}{192}\zeta+\frac{9}{64}\zeta^{3}\,$,
where
$\zeta:=2\left[\frac{\Gamma(\frac{3}{4})}{\Gamma(\frac{1}{2})}\right]^{4}=2\frac{\boldsymbol{\mathsf{E}}\left(\frac{1}{\sqrt{2}}\right)}{\boldsymbol{\mathsf{K}}\left(\frac{1}{\sqrt{2}}\right)}-1\,.$

### v2

@@ 7. Discussion and concluding remarks
For the case $d=2$, the analysis involves a spectral problem that can be mapped, by allowing a certain parameter ratio to assume imaginary values, to that studied by Chetrite et al. [2] for the Kraichnan ensemble of velocity fields. We computed, as they did, the first two cumulants in terms of elliptic functions. We also showed how these calculations relate, through the work of [4], to the one-dimensional Anderson model at zero energy [17, 10, 24], and to the one-dimensional Dirac equation with a random mass [23].
For $d\geq 2$, we embedded the problem in a $k$-dependent family such that, when $k=0$, the problem is solvable. Treating $k$ as a perturbation parameter, we then obtained expansions in ascending powers of $k^{2}$ for a function $\Lambda(k,\ell)$ that, suitably rescaled, coincides in the continuum limit with $L(\ell)$ when $k^{2}=1/d$. The first few terms of this expansion were computed explicitly for $d\in\{2,3\}$. We observed that the radius of convergence relates to the spectral gap separating the dominant eigenvalue of the transfer operator from the rest of its spectrum, and that it decreases with $\ell$ so that, when $k^{2}=1/d$, the expansions only converge for moderate values of $\ell$. Nevertheless, they provided an effective means of computing the cumulants, which depend only on the local behaviour of the generalised Lyapunov exponent at $\ell=0$.
We also showed that, for every $0\leq k^{2}\leq 1/d$, $\ell\mapsto\Lambda(k,\ell)$ is— still in the continuum limit and suitably scaled— the generalised Lyapunov exponent of a product of matrices associated with the renewing flow obtained by inserting time intervals of pure strain; the number $1/d-k^{2}$ then measures the strength of the disorder in the pure strain, relative to that in the pure-strain-free version. Whether there is, for the case $d>2$, a simple mapping between this model and the Kraichnan model studied by Chetrite et al. is an interesting open question.
@@ 7.1. Comparison with the calculations of Haynes and Vanneste
As mentioned in the introduction, Haynes & Vanneste’s objective is to determine conditions under which the decay rate of the covariance $\Gamma(0,t)$ of a passive scalar, evolving in a spatially periodic domain, may be deduced from the stretching characteristics of the random flow. The main conclusion of the analysis presented in their section III, devoted to the continuum limit, is that, as long as the period of the spatial domain is a small enough integer multiple of that of the velocity field, then
$\lim_{t\rightarrow\infty}\frac{\ln\Gamma({\mathbf{x}},t)}{t}=\Lambda(0)-\frac{2\pi^{2}\Lambda^{\prime\prime}(0)}{\ln^{2}\kappa}+o\left(\frac{1}{\ln^{2}\kappa}\right)\;\;\text{as $\kappa\rightarrow 0+$}$
where $\Lambda(0)$ and $\Lambda^{\prime\prime}(0)$ are expressible in terms of what we denote in the present paper $\Lambda(k,\ell)$. This is their Equation (3.8), reproduced here using their notation, which uses $\kappa$ instead of $D$ for the diffusivity constant. In the statement just above their equation (3.21), they also give the numerical values $\Lambda(0)\approx 0.226$ and $\Lambda^{\prime\prime}(0)\approx-0.45$ but, when one compares their equations (3.8) and (3.21), it becomes clear that $\Lambda^{\prime\prime}(0)\approx-0.45$ was intended.
The precise correspondence between their function $\Lambda$ and our own $\Lambda(k,\ell)$ is clouded by some unfortunate clashes of notation. To clarify it, let us take as starting point Haynes & Vanneste’s Equation (3.19), obtained after separating the spatial variable ${\mathbf{x}}=(x,y)$ from the temporal variable $t$:
@@
(7.1) ${\mathbf{x}}=(x,y)\sim x=(x_{1},x_{2})\,,\;\;\kappa\sim D\,,\;\;\sigma\sim\ell+1\,,\\ f_{\sigma}(\theta)\sim v(\theta)\,,\;\;\Lambda(\sigma)\sim-\Lambda\left(k,\ell\right)\Bigl|_{k=\frac{1}{\sqrt{2}}}\,.$
In particular, we see that, in the work of Haynes and Vanneste, it is the value taken by $\Lambda(1/\sqrt{d},\ell)$ at its minimum
(7.2) $\ell=\ell_{\min}:=-\frac{d}{2}\,$,
as well as the value of the second derivative there, that govern the temporal decay of the covariance.
It turns out that $\ell_{\min}$ is small enough to ensure that the partial sums of the series for $\Lambda(k,\ell)$ in powers of $k^{2}$ converge nicely at $k^{2}=1/d$ when $\ell=\ell_{\min}$. In particular, for $d=2$, we estimate— on the basis of the first 64 terms computed with the help of the mathematical software Maple— that, for $\ell=\ell_{\min}=-1$, the radius of convergence is approximately $1$. This expansion translates into
(7.3) $\Lambda\left(k,\ell_{\min}\right)=-\frac{1}{2}+\frac{1}{2}k^{2}+\frac{9}{128}k^{4}+\frac{9}{256}k^{6}+\frac{47241}{2097152}k^{8}\\ +\frac{67995}{4194304}k^{10}+\frac{13407669}{1073741824}k^{12}+\frac{21598857}{2147483648}k^{14}+\cdots$
and
(7.4) $\frac{\partial^{2}\Lambda}{\partial\ell^{2}}\left(k,\ell_{\min}\right)=1-k^{2}-\frac{5}{32}k^{4}-\frac{5}{64}k^{6}-\frac{13109}{262144}k^{8}\\ -\frac{18847}{524288}k^{10}-\frac{7424679}{268435456}k^{12}-\frac{11948355}{536870912}k^{14}+\cdots$
For $k^{2}=1/2$, the partial sums converge well and, by comparing them, we deduce the following values, believed to be correct to the number of decimals displayed:
$\Lambda\left(k,\ell_{\min}\right)\Bigl|_{k=\frac{1}{\sqrt{2}}}=-0.2257817708\ldots\;\;\text{and}\;\;\frac{\partial^{2}\Lambda}{\partial\ell^{2}}\left(k,\ell_{\min}\right)\Bigl|_{k=\frac{1}{\sqrt{2}}}=0.4461902388\ldots$
7.2. The spectrum of the transfer operator and its dependence on $\tau$, $k$ and $\ell$
Equation (1.30) and Proposition 1.1 imply that the eigenvalues of the transfer operator, in the continuum limit, are of the form
$\lambda=\lambda_{l}:=1-\tau^{2}l(l+d-2)\,,\;\;\text{for some}\;l\in\{0,1,2,\ldots\},$
to leading order as $\ell$ and $k$ both tend to zero. In particular, the leading eigenvalue appearing in Equation (1.25) is indeed strictly positive in this limiting case. What becomes of the spectrum as the parameters $k$, $\ell$ and $\tau$ vary is an open question; a systematic numerical study of the convergence properties of the expansions in powers of $k$ obtained in this paper could provide a useful starting point for investigating it.
@@ 7.4. Beyond the continuum limit
where, according to Equation (1.12), $\alpha=2\tau\cos\theta$ and $\beta=2\tau\cos\theta^{\prime}$, with $\theta$ and $\theta^{\prime}$ drawn independently from the uniform distribution in $(-\pi,\pi)$.
Let us take the calculation of the first cumulant $c_{1}$ in Equation (1.18) as a modest objective. The following result, whose proof we omit, is a simple application of Furstenberg’s formula [13] for the Lyapunov exponent of a product of independent identically-distributed random matrices:
@@
${c}_{1}=\int_{\mathbb{R}}\nu(dz)\,{\mathbb{E}}\left(\ln\left|\frac{\alpha+z}{z}\right|\right)\,.$
@@
${c}_{1}=-2\int_{\mathbb{R}}\nu(dz)\,\ln|z|\,.$
The problem of computing $c_{1}$ therefore reduces to that of finding the distribution of the random continued fraction implied by Equation (7.8). This continued fraction has the same form as that considered by Dyson in his study of a random linear chain [11, 12]. There is, however, one significant novelty: in Dyson’s case the $\alpha_{j}$ and $\beta_{j}$ represent random masses and Hooke constants, and so assume strictly positive values; by contrast, in the present case, the $\alpha_{j}$ and $\beta_{j}$ can assume negative values. This is problematic (and the situation gets worse as $\tau$ increases) because Equation (7.8) implies that the probability of dividing by a number in an interval centered on zero is proportional to the size of the interval.
@@
${c}_{1,0}=0\,,\;\;{c}_{1,1}=\frac{1}{\pi}G^{-2}\,,\;\;{c}_{1,2}=\frac{1}{24}+\frac{3}{8\pi^{2}}G^{-4}\,,\;\;{c}_{1,3}=\frac{5}{192\pi}G^{-2}+\frac{9}{64\pi^{3}}G^{-6}\,$,
where $G$ is the Gauss lemniscate constant; see Equation (4.37).